OpenStudy (anonymous):

There are 400 types of poisonous snakes in the world. Given that this number represents 13 1/3% of the total number of snake varieties, how many varieties are there in the world?

OpenStudy (anonymous):

Well you could do proportions \[\frac{ 400 }{ x }=\frac{ 13.333 }{ 100 }\]Then cross multiply \[40000=13\frac{ 1 }{ 3 }x\]Then divide by 13 1/3 to get x or total varieties of species \[x=3000\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your Welcome and Good Luck :)